OpenStudy (anonymous):

A 3-mile cab ride costs $5.75. A 5-mile cab ride costs $8.25. Part1: Find a linear equation that models the cost c as a function of the distance d. (Hint: Create a table with data points and use the data from the table to write the equation.) Part 2: How much will a 10-mile cab ride cost?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (aripotta):

so let's make some points. 3 miles costs $5.75, so (3, 5.75). 5 miles is $8.25 so (5, 8.25) now we can find the slope \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

How would you figure out what it cost for each mile?

OpenStudy (aripotta):

that's what the slope is

OpenStudy (aripotta):

\[\frac{ 8.25-5.75 }{ 5-3 }\]

OpenStudy (michele_laino):

please note that, using your data and conjecturing a linear relationship, namely if I caal with c the cost and with x number of miles, then I conjecture this: c=a*x+b

OpenStudy (anonymous):

\[\frac{ 2.50 }{ 2 }\]

OpenStudy (michele_laino):

inserting your numerical data, I can write: \[5.75=3a+b\] \[8.75=5a+b\] solving those above equation in order to find a and b, we have: a=1.25, and b=4.00

OpenStudy (michele_laino):

so the function c, is, please continue, namely insert values for a and b into the above equation

OpenStudy (anonymous):

So the equation would be C = 1.25d + 4.00 ?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

On the answer key it says: C = 1.25d + 2.00

OpenStudy (anonymous):

I was trying to figure out how the answer would be C = 1.25d + 2.00

OpenStudy (aripotta):

2.5/2 is 1.25, so that's the slope. once we put that and one of the points, we put it in point-slope form to get y - 5.75 = 1.25(x - 3)

OpenStudy (michele_laino):

@Answers101 please note that it has been necessary to conjecture a linear relationship between number of miles and cost

OpenStudy (aripotta):

once we convert it, we get y - 1.25x + 2

OpenStudy (aripotta):

y =*

OpenStudy (anonymous):

So the answer is C = 1.25d + 2.00 because we had to convert the linear relationship between the number of miles & the cost. ?

OpenStudy (michele_laino):

@Answers101 sorry you are right: c=1.25 x+2.00

OpenStudy (michele_laino):

@Answers101 please substitute x=10, and then you will get your answer